# Wider or Deeper? Scaling LLM Inference-Time Compute with Adaptive Branching Tree Search

### v3

@@ Abstract
Recent advances demonstrate that increasing inference-time computation can significantly boost the reasoning capabilities of large language models (LLMs). Although repeated sampling (i.e., generating multiple candidate outputs) is a highly effective strategy, it does not leverage external feedback signals for refinement, which are often available in real tasks like coding. In this work, we propose Adaptive Branching Monte Carlo Tree Search (AB-MCTS), a novel inference-time framework that generalizes repeated sampling with principled multi-turn exploration and exploitation. At each node in the search tree, AB-MCTS dynamically decides whether to “go wider” by expanding new candidate responses or “go deeper” by revisiting existing ones based on external feedback signals. We evaluate our method on complex coding and engineering tasks using frontier models. Empirical results show that AB-MCTS outperforms both repeated sampling and standard MCTS, underscoring the importance of combining the response diversity of LLMs with multi-turn solution refinement for effective inference-time scaling.
@@ 1 Introduction
Several approaches [6, 7, 9, 10, 11, 15] have been proposed for exploration and exploitation in such multi-turn settings, but the majority were designed before the power of inference-time scaling was fully recognized. Consequently, these methods use a fixed “width”, i.e., they treat the number of answers generated from a single prompt as a fixed hyperparameter. For example, methods based on standard Monte Carlo Tree Search (MCTS) use a fixed branching factor (i.e., the number of child nodes per state) as a hyperparameter [9, 10, 11, 15]. As demonstrated by the success of repeated sampling, effective inference-time scaling requires leveraging a diverse and vast output space, so a fixed width hinders scaling.
@@ 2 Related Work
Inference-Time Scaling by Multiple Answer Generation. Since the community has come to appreciate the power of inference-time scaling, the strategy that has been studied widely is repeated sampling, in which the model generates many candidate answers and selects the best one [1, 2, 3, 35]. Although empirically strong and widely used, repeated sampling leaves obvious room for improvement because it does not refine its candidates using external feedback [4, 5]. Before the era of large-scale inference-time compute, a variety of task-specific strategies were proposed for relatively small scales; examples include tree expansions directed by LLMs [6] and Bayesian methods [7]. The most common alternative, however, relies on MCTS. LATS [9], RAP [15], SWE-Search [10], and RepoUnderstander [11] all employ MCTS in different settings while leaving the core algorithm unchanged; in every case, they use the conventional version with a fixed branching factor. Because we focus on the core search algorithms, we group these previous works under the term “standard MCTS” and compare them with our proposed framework in the experiments.
Progressive Widening in MCTS. Progressive widening [36, 37] is a classic technique that gradually increases actions considered per node. It was designed for games with unique actions and no side information for untried moves, relying on visit-count heuristics. Our approach differs as new branches are sampled from the same LLM. This homogeneity in generation allows a principled statistical rule for choosing between widening and deepening.
@@ 3.2 Adaptive Branching MCTS
In standard MCTS, only leaf nodes are selected and expanded, (i.e. each node is expanded at most once), and the expansion adds a fixed number of child nodes. However, since each query to an LLM at non-zero temperature can yield different outputs from the same prompt, the branching factor is theoretically infinite. To accommodate such unbounded branching, we relax the standard MCTS constraints and allow selection and expansion of non-leaf nodes. Moreover, recent studies [3] suggest that drawing many outputs from the same prompt at non-zero temperature can improve performance. Allowing unbounded branching enables us to fully exploit these varied samples, whereas restricting the branching factor could miss correct answer generations and undermine overall performance.
Adaptive Branching via the GEN Node. To fully leverage the potential performance improvement from unbounded branching, we allow nodes that have already been expanded once to be expanded again and further branched, unlike in standard MCTS. To explicitly represent the action of generating new child nodes, we introduce a GEN node. Every node $N$ (including newly expanded ones during iterations) has a GEN node as a child. When the GEN node with parent node $N$ is selected during the selection step, we expand $N$ to add a new child to it. Algorithm 1 outlines this approach, called Adaptive Branching Monte Carlo Tree Search (AB-MCTS).
@@
[…]
9:$N\leftarrow N_{\text{next}}$EndWhile
[…]
@@
Thompson Sampling for Node Selection. In our proposed methods, we employ a Bayesian approach with Thompson sampling for node selection. In standard MCTS, the usual selection policy is based on UCT score [38], but since GEN nodes have no child nodes, UCT scores cannot be assigned. Thompson sampling not only resolves this issue but also enables parallel node expansion. This is particularly beneficial when evaluating node scores is time-consuming, as in the case of MLE-Bench (See Appendix B.1 for MLE-Bench details). Next, we discuss how to utilize Thompson sampling to choose between a GEN node and existing child nodes during the selection step at node $N$. At node $N$, we have $1+n_{\rm child}$ possible actions: choosing a GEN node or one of the $n_{\rm child}$ existing nodes. Since our purpose is to expand the node with score as high as possible, we model the probability distributions of the score of an eventually expanded node at that iteration after choosing actions (i.e. the score of the node $N_{\rm new}$ at Line 5 of Algorithm 1). Thompson sampling proceeds by first sampling scores from the modeled probability distributions for all the actions, and then choosing the action with the highest sampled score. A key question is how to model the expanded node’s score probability distribution for each action. We address this with two strategies: a mixed Bayesian model (AB-MCTS-M) and a node aggregation method (AB-MCTS-A). In both cases, we model the score probability distributions by Bayesian posterior predictives, but with different statistical models.
@@ 3.3 AB-MCTS-M: Adaptive Branching MCTS with Mixed Model
During the selection at node $N$, estimating the score of a GEN node is challenging because it does not directly correspond to a specific LLM-generated observation. To address this, we employ a separate, node-specific mixed model fitted individually at each node during the selection step. In a mixed model, we assign a “group” label to each node, and nodes sharing the same label share a model parameter. This shared model parameter captures common characteristics of each group, such as the answer quality of the parent node of each group. In our setting, denoting each child node of $N$ as $N_{j}$ ($j=1,\dots,n_{\rm child}$), each group corresponds to the set of nodes $T_{\rm sub}(N_{j})$ (i.e. all the nodes in the subtree under $N_{j}$, see Figure 2), which consist of answers generated by refinement from the answer of the common ancestor node $N_{j}$. Our assumption is that even after multiple stages of refinements, the quality associated with $N_{j}$’s answer continues to be represented by this shared parameter (see Appendix A.2.2 for model parameter details). In contrast, the GEN node corresponds to a new group with no scores observed yet, but we note that, since it uses the shared parameter as the other groups, the other groups’ node scores affect the GEN node’s score distribution as well.
Algorithm Outline. As described in Section 3.2, our goal is to model the probability distribution of the scores for the nodes that will eventually be expanded after choosing a GEN node or any child node during the selection step at node $N$. AB-MCTS-M assigns each subtree under $N_{j}$, denoted as $T_{\text{sub}}(N_{j})$, as a distinct group (see Figure 2 for example subtree). The mixed model then leverages observed scores from these groups to compute the posterior predictive distributions of expected scores for new nodes generated from each group (See Figure 2 for a schematic illustration and Appendix A.2.2 for a detailed formulation of the mixed model.) We sample the scores from all the groups (the GEN node and $T_{\text{sub}}(N_{j})$) using calculated posterior predictives. If the GEN node’s sampled score is highest, we call $f_{\rm LLM}$ to generate a new child node. Otherwise, we choose the child node $N_{j}$ with the highest score and continue the sampling step.
Score Backup Mechanism. When a new node $N$ is created, its observed score is added to the histories of $N$ and its ancestors. This cumulative record is used to update the posterior distributions in the mixed model. The observed score is not backed up to a GEN node, but it indirectly influences the GEN node’s score probability distribution through the shared parameters in the mixed model (see Appendix A.2.2 for details).
@@ 3.4 AB-MCTS-A: Adaptive Branching MCTS with Node Aggregation
For probability distributions, we use exponential families with conjugate priors to enable analytic and efficient posterior updates. We employ two variants: (1) AB-MCTS-A (Gaussian), using a normal-inverse-$\chi^{2}$ prior for unbounded scores, and (2) AB-MCTS-A (Beta), using a Beta prior for scores in $[0,1]$. See Figure 3 for example tree, and Appendix A.3 for detailed parameter updates.
Score Backup Mechanism. During score-backup operations, the expanded node score is backed up to the GEN node, which led to the expansion of that node and the GEN node’s ancestors. As we can see from Figure 3, a GEN node’s ancestors include only generated nodes and CONT nodes, so the score is backed up to a GEN node only from the node that is created by choosing that GEN node. The backed-up score is used to update prior probability distribution parameters.
@@ 4.1 Experimental Setup
Models. We perform our experiments using GPT-4o (gpt-4o-2024-08-06[18], and DeepSeek-V3 (deepseek-chat[19]. Each LLM generates a complete solution in a single API call. We define the generation budget as the maximum number of API calls and set the generation budget, which we set to $1282^{7}=$ 128. The temperature was set to 0.6 for GPT-4o following [3] and 1.0 for DeepSeek-V3 following the official document.
Baselines. We benchmark AB-MCTS against three representative approaches. (1) Repeated Sampling (Best-of-$n$) [1, 3, 39] independently generates up to $n$ candidate solutions from a single LLM prompt, a simple yet competitive baseline for coding tasks. (2) Sequential Refinement [4] iteratively improves each solution by re-prompting the LLM with its own output and feedback. (3) Standard MCTS backs projects such as LATS [9], RAP [15], SWE-Search [10], and RepoUnderstander [11]. As we design a core search algorithm, we group and compare them as standard MCTS. Following the previous work [9, 15], each expansion adds five child nodes, and the search proceeds until it reaches the $2^{7}$ nodes. Hyper-parameters for AB-MCTS are summarized in Appendix B.2.
Table 1: Performance of AB-MCTS against baselines across benchmarks and models. This table compares AB-MCTS with the baseline methods. Evaluations were performed on LiveCodeBench, CodeContest, and ARC-AGI using GPT-4o and DeepSeek-V3 with a maximum generation budget ($2^{7}$). Each entry provides a performance score (higher values are better) and its corresponding rank (in parentheses, 1st is best). The "Avg. Rank" column shows the average rank across all settings.
@@
Repeated Sampling [1, 3, 39] 37.8 (4) 40.7 (6) 37.9 (4) 43.2 (5) 15.0 (1) 18.6 (1) 3.5
Sequential Refinement [4] 37.8 (4) 41.6 (5) 30.1 (6) 41.6 (6) 8.7 (6) 10.0 (6) 5.5
Standard MCTS [9, 15] 36.7 (6) 43.2 (1) 37.5 (5) 43.8 (3) 9.0 (5) 14.0 (5) 4.2
AB-MCTS-M 38.9 (2) 43.0 (2) 40.6 (1) 44.6 (2) 12.3 (4) 16.0 (3) 2.3
AB-MCTS-A (Gaussian) 39.1 (1) 42.5 (3) 40.2 (3) 43.4 (4) 13.0 (3) 18.3 (2) 2.7
AB-MCTS-A (Beta) 38.7 (3) 42.3 (4) 40.4 (2) 44.8 (1) 14.0 (2) 16.6 (4) 2.7

### v5

@@ Abstract
Recent advances demonstrate that increasing inference-time computation can significantly boost the reasoning capabilities of large language models (LLMs). Although repeated sampling (i.e., generating multiple candidate outputs) is a highly effective strategy, it does not leverage external feedback signals for refinement, which are often available in real tasks like coding. In this work, we propose Adaptive Branching Monte Carlo Tree Search (AB-MCTS), a novel inference-time framework that generalizes repeated sampling with principled multi-turn exploration and exploitation. At each node in the search tree, AB-MCTS dynamically decides whether to “go wider” by expanding new candidate responses or “go deeper” by revisiting existing ones based on external feedback signals. We evaluate our method on complex coding and engineering tasks using frontier models. Empirical results show that AB-MCTS outperforms both repeated sampling and standard MCTS, underscoring the importance of combining the response diversity of LLMs with multi-turn solution refinement for effective inference-time scaling. Code is available at: https://github.com/SakanaAI/treequest.
@@ 1 Introduction
Several approaches [6, 7, 9, 10, 11, 15] have been proposed for exploration and exploitation in such multi-turn settings, but the majority were designed before the power of inference-time scaling was fully recognized. Consequently, these methods use a fixed “width”, i.e., they treat the number of answers generated from a single prompt as a fixed hyperparameter. For example, methods based on standard Monte Carlo Tree Search (MCTS) use a fixed branching factor (i.e., the number of child nodes per state) as a hyperparameter [9, 10, 11, 15]. As demonstrated by the success of repeated sampling, effective inference-time scaling requires leveraging a diverse and vast output space, thus, providing substantial evidence that a fixed width hinders scaling.
@@ 2 Related Work
Inference-Time Scaling by Multiple Answer Generation. Since the community has come to appreciate the power of inference-time scaling, the strategy that has been studied widely is repeated sampling, in which the model generates many candidate answers and selects the best one [1, 2, 3, 35]. Although empirically strong and widely used, repeated sampling leaves obvious room for improvement because it does not refine its candidates using external feedback [4, 5]. Before the era of large-scale inference-time compute, a variety of task-specific strategies were proposed for relatively small scales; examples include tree expansions directed by LLMs [6] and Bayesian methods [7]. LATS [9], RAP [15], SWE-Search [10], and RepoUnderstander [11] combine LLMs with MCTS, primarily targeting sequential decision making. In this context, nodes represent states and edges represent actions, which may involve interaction with an environment. LATS utilizes API calls and code execution as actions to solve tasks. RAP addresses the process of solving block-moving puzzles and mathematical word problems step-by-step. SWE-Search explores sequences of actions such as searching, editing, and running tests to resolve issues within a software repository. RepoUnderstander employs MCTS for exploration on a repository knowledge graph. The application of LATS to coding tasks [9, Section 5.2] aligns with the context of multiple-answer generation in this paper and corresponds to what we refer to as “standard MCTS” in our experiments.
Progressive Widening in MCTS. Progressive widening (PW) [36, 37] is a classic technique that gradually increases actions considered per node. It was designed for games with unique actions and no side information for untried moves, relying on visit-count heuristics. Complementary to PW, Sokota et al. [38] propose “abstraction refining”, which groups similar successors using a decreasing similarity threshold and shows advantages over PW under equal simulation budgets in stochastic domains. Our approach differs as new branches are sampled from the same LLM. This homogeneity in generation allows a principled statistical rule for choosing between widening and deepening.
@@ 3.2 Adaptive Branching MCTS
In standard MCTS, only leaf nodes are selected and expanded, (i.e., each node is expanded at most once), and the expansion adds a fixed number of child nodes. However, since each query to an LLM at non-zero temperature can yield different outputs from the same prompt, the branching factor is theoretically infinite. To accommodate such unbounded branching, we relax the standard MCTS constraints and allow selection and expansion of non-leaf nodes. Moreover, recent studies [3] suggest that drawing many outputs from the same prompt at non-zero temperature can improve performance. Allowing unbounded branching enables us to fully exploit these varied samples, whereas restricting the branching factor could miss correct answer generations and undermine overall performance.
Adaptive Branching via the GEN Node. To fully leverage the potential performance improvement from unbounded branching, we allow nodes that have already been expanded once to be expanded again and further branched, unlike in standard MCTS. To explicitly represent the action of generating new child nodes, we introduce a GEN node. Every node $N$ (including newly expanded ones during iterations) has a GEN node as a child. When the GEN node with parent node $N$ is selected during the selection step, we expand $N$ by adding a new child node. Algorithm 1 outlines this approach, called Adaptive Branching Monte Carlo Tree Search (AB-MCTS).
@@
1: function AB-MCTS($n_{\rm nodes}$)
2:   $T\leftarrow$ InitializeTree( )
3:   for $n=1,\dots,n_{\rm nodes}$ do
4:     $N\leftarrow$ SelectExpansionTarget($T$) $\triangleright$ Step 1. Select an expansion target
5:     $N_{\text{new}}\leftarrow$ Expand(NN, $T$) $\triangleright$ Step 2. Expand the selected node to generate a child
6:     ScoreBackUp($N_{\text{new}}$, $T$) $\triangleright$ Step 3. Backup the score from the generated node
7:   return SelectBest($T$)
8: function SelectExpansionTarget($T$)
9:   $N\leftarrow$ GetRoot($T$)
10:   while not IsLeaf(NN) do
11:     $N_{\text{next}}\leftarrow$ SelectChild(NN, $T$) $\triangleright$ Detailed in Sections 3.3 and 3.4
12:     if IsGenNode($N_{\text{next}}$) then $\triangleright$ If a GEN node is selected, branch off from the node
13:      break
14:     $N\leftarrow N_{\text{next}}$
15:   return NN
@@
The UCT score is inapplicable to our AB-MCTS because GEN nodes make the problem fundamentally different from a standard multi-armed bandit problem, for which UCT was designed. In standard MCTS, the arms (branches) are static. In contrast, the GEN node in AB-MCTS dynamically generates new arms. This special problem setting, where arms are generated on the fly, prevents the direct application of UCT. We, therefore, adopt a Bayesian probabilistic model. This enables Thompson sampling based on the posterior distribution and obviates the need for complex UCB-style confidence bound analysis.
Thompson Sampling for Node Selection. In our proposed methods, we employ a Bayesian approach with Thompson sampling for node selection. Here, we employ Thompson sampling because GEN nodes do not have child nodes, making it impossible to compute their UCT scores. In addition, Thompson sampling has the advantage of allowing node expansion in parallel. This is particularly beneficial when evaluating node scores is time-consuming, as in the case of MLE-Bench (See Appendix B.1 for MLE-Bench details).
Concretely, during SelectChild step at line 11 of Algorithm 1, we employ Thompson sampling to decide between expanding a GEN node or selecting from existing child nodes at node $N$. Let $N$ be a node with potential actions
$A_{N}=\{a_{0},a_{1},\dots,a_{n_{\rm child}}\},$
where the action $a_{0}$ corresponds to choosing the GEN node, and $a_{1},\dots,a_{n_{\rm child}}$ correspond to choosing the already-existing child nodes. Suppose $P_{N}(r\mid a_{i})$ is the posterior predictive distribution of the score $r$ for an eventually expanded new node ($N_{\text{new}}$ at line 5 of Algorithm 1) if we choose the action $a_{i}$ at node $N$. Then Thompson sampling proceeds by,
  1. 1.
    Calculate $P_{N}(r\mid a_{j})$ for each action $a_{j}$ at node $N$.
  2. 2.
    Draw scores $r_{N_{\text{new}},a_{j}}$ from $P_{N}(r\mid a_{j})$ for each action $a_{j}$.
  3. 3.
    Select $\hat{a}=\arg\max_{a_{j}\in A_{N}}r_{N_{\text{new}},a_{j}}$.
This three-step process corresponds to a single call to SelectChild.
A key question is how to perform step 1, i.e., how to model and calculate $P_{N}(r\mid a_{j})$ for all $a_{j}$, in particular for $j=0$ (i.e., GEN node). We address this with two strategies: a mixed Bayesian model (AB-MCTS-M) and a node aggregation method (AB-MCTS-A). In both cases, we model the score probability distributions by Bayesian posterior predictives, but with different statistical models.
@@ 3.3 AB-MCTS-M: Adaptive Branching MCTS with Mixed Model
To model $P_{N}(r\mid a_{j})$, we employ a node-specific mixed model fitted individually at each node $N$. That is, we fit a separate model for each $N$ every time SelectChild in Algorithm 1 is invoked. Denoting $r_{N_{\text{new}},a_{j}}\sim P_{N}(r\mid a_{j})$ as a score of an eventually expanded node $N_{\text{new}}$ if we choose an action $a_{j}$ at $N$, our mixed model is given as:
$\begin{gathered}r_{N_{\text{new}},a_{j}}=\alpha_{j}+\sigma_{y}\epsilon_{N_{\text{new}}},\quad\alpha_{j}=\mu_{\alpha}+\sigma_{\alpha}\epsilon_{j},\\ \epsilon_{N_{\text{new}}}\sim\mathcal{N}(0,1),\quad\epsilon_{j}\sim\mathcal{N}(0,1),\end{gathered}$ (1)
Here, $\alpha_{j}$ is a “group-level” intercept capturing the quality of the base solution at $N_{j}$, while $\sigma_{y}\epsilon_{N_{\text{new}}}$ represents per-instance noise. To fit this model, we place priors on the hyperparameters ($\mu_{\alpha}$, $\sigma_{\alpha},\sigma_{y}$) and employ Markov Chain Monte Carlo (MCMC) to sample from their posterior distribution. The GEN node (action $a_{0}$) is treated as a newly introduced group without its own direct observations. However, its group-level intercept $\alpha_{0}$ is inferred not from the prior alone but rather from the posterior distribution over $\mu_{\alpha}$ and $\sigma_{\alpha}$, which is informed by the other observed data. We assume that even after multiple refinement stages, the quality associated with the answer at node $N_{j}$ continues to be captured by this shared parameter (see Appendix A.3.2 for further details).
Algorithm Outline. To model $P_{N}(r\mid a_{j})$, AB-MCTS-M assigns each subtree under $N_{j}$, denoted as $T_{\text{sub}}(N_{j})$, as a distinct group $j$ (see Figure 2 for example subtree). The mixed model leverages observed scores from these groups to compute the posterior predictive distributions of expected scores for new nodes generated from each group (See Figure 2 for a schematic illustration). We sample the scores from all the groups (the GEN node and $T_{\text{sub}}(N_{j})$) using calculated posterior predictives. If the GEN node’s sampled score is highest, we call $f_{\text{LLM}}$ to generate a new child node. Otherwise, we choose the child node $N_{j}$ with the highest score and continue the sampling step.
Score Backup Mechanism. When a new node $N$ is created, its observed score is added to the histories of $N$ and its ancestors. This cumulative record is used to update the posterior distributions in the mixed model. The observed score is not backed up to a GEN node, but it indirectly influences the GEN node’s score probability distribution through the shared parameters in the mixed model (see Appendix A.5 for a detailed walkthrough).
@@ 3.4 AB-MCTS-A: Adaptive Branching MCTS with Node Aggregation
To model $P_{N}(r\mid a_{j})$, we utilize exponential family distributions with conjugate priors, enabling analytical and efficient posterior updates. We employ two variants:
  1. 1.
    AB-MCTS-A (Gaussian), using a normal-inverse-$\chi^{2}$ prior for unbounded scores: $P_{N}(r\mid a_{j})=p(r\mid\{r_{k}\}_{k=1}^{K})=\mathcal{N}(r\mid\hat{m},\tfrac{\sigma^{2}}{\hat{\kappa}})\chi^{-2}(\sigma^{2}\mid\hat{\nu},\hat{\tau}^{2})$, and
  2. 2.
[…]
where $r_{k}$ represents the scores backed up to the node $N_{j}$ (GEN node, CONT node or LLM-generated child nodes; see Figure 3 for example tree), where $\hat{m},\hat{\kappa},\hat{\nu},\hat{\tau},\hat{\alpha},\hat{\beta}$ are determined from observed scores $r_{k}$ and updated as these scores are backed up. The detailed parameter update rules are given in Appendix A.4.
Score Backup Mechanism. During score-backup operations, the expanded node score is backed up to the GEN node which led to the expansion of that node and the GEN node’s ancestors (see Appendix A.5 for a detailed walkthrough). As we can see from Figure 3, a GEN node’s ancestors include only generated nodes and CONT nodes, so the score is backed up to a GEN node only from the node that is created by choosing that GEN node. The backed-up score is used to update prior probability distribution parameters.
@@ 4.1 Experimental Setup
Models. We perform our experiments using GPT-4o (gpt-4o-2024-08-06[18], and DeepSeek-V3 (deepseek-chat[19]. Each LLM generates a complete solution in a single API call. We define the generation budget as the maximum number of API calls and set it to $2^{7}=128$. The temperature was set to 0.6 for GPT-4o following [3] and 1.0 for DeepSeek-V3 following the official documentation.
Baselines. We benchmark AB-MCTS against three representative approaches. (1) Repeated Sampling (Best-of-$n$) [1, 3, 39] independently generates up to $n$ candidate solutions from a single LLM prompt, a simple yet competitive baseline for coding tasks. (2) Sequential Refinement [4] iteratively improves each solution by re-prompting the LLM with its own output and feedback. (3) Standard MCTS follows the configuration from LATS [9, Section 5.2] (See also Appendix C.8). Each expansion adds five child nodes, and the search proceeds until it reaches the $2^{7}$ nodes, with the final expansion creating only three nodes to meet this limit precisely. Hyper-parameters for AB-MCTS are summarized in Appendix B.2.
Table 1: Performance of AB-MCTS against baselines across benchmarks and models. This table compares AB-MCTS with the baseline methods. Evaluations were performed on LiveCodeBench, CodeContest, and ARC-AGI using GPT-4o and DeepSeek-V3 with a maximum generation budget ($2^{7}$). Each entry provides a performance score (higher values are better) and its corresponding rank (in parentheses, 1st is best). The “Avg. Rank” column shows the average rank across all settings.
@@
Repeated Sampling 37.8 $\pm$ 0.5 (4) 40.7 $\pm$ 1.9 (6) 37.9 $\pm$ 0.3 (4) 43.2 $\pm$ 0.9 (5) 15.0 $\pm$ 1.0 (1) 18.6 $\pm$ 1.0 (1) 3.5
Sequential Refinement 37.8 $\pm$ 2.4 (4) 41.6 $\pm$ 0.6 (5) 30.1 $\pm$ 0.3 (6) 41.6 $\pm$ 0.9 (6) 8.7 $\pm$ 0.9 (6) 10.0 $\pm$ 0.6 (6) 5.5
Standard MCTS 36.7 $\pm$ 1.0 (6) 43.2 $\pm$ 2.1 (1) 37.5 $\pm$ 0.0 (5) 43.8 $\pm$ 0.9 (3) 9.0 $\pm$ 1.5 (5) 14.0 $\pm$ 1.5 (5) 4.2
AB-MCTS-M 38.9 $\pm$ 1.9 (2) 43.0 $\pm$ 1.5 (2) 40.6 $\pm$ 1.0 (1) 44.6 $\pm$ 0.9 (2) 12.3 $\pm$ 1.2 (4) 16.0 $\pm$ 1.0 (3) 2.3
AB-MCTS-A (Gaussian) 39.1 $\pm$ 1.9 (1) 42.5 $\pm$ 1.5 (3) 40.2 $\pm$ 1.7 (3) 43.4 $\pm$ 0.9 (4) 13.0 $\pm$ 3.6 (3) 18.3 $\pm$ 0.6 (2) 2.7
AB-MCTS-A (Beta) 38.7 $\pm$ 1.2 (3) 42.3 $\pm$ 0.8 (4) 40.4 $\pm$ 0.3 (2) 44.8 $\pm$ 0.6 (1) 14.0 $\pm$ 2.1 (2) 16.6 $\pm$ 0.6 (4) 2.7